# Can LLMs Handle WebShell Detection? Overcoming Detection Challenges with Behavioral Function-Aware Framework

### v4

@@ 2.4 LLM-Based Detection Framework
We integrate the BFAD framework into the LLM-based detection system, which combines the context-aware code extraction strategy and WBFP to optimize the use of LLMs for WebShell detection. By incorporating both critical code regions and global context, along with behaviorally relevant demonstrations, our framework enhances the ability of the LLM to accurately identify malicious code patterns.
The input to the LLM consists of two main components: (a) a system directive that defines the model’s role as a cybersecurity expert, and (b) a user query that contains the extracted critical code segments and a selected ICL demonstration. To balance efficiency and performance, we limit the user query to one ICL demonstration, which reduces computational overhead while preserving sufficient context for reliable detection. Our prompt is described in detail in Appendix B.
@@ 3.1 Dataset Overview
We constructed a comprehensive dataset consisting of 21,665 benign PHP programs and 4,929 WebShells. The benign programs were obtained from established open-source PHP projects to ensure applicability to real-world scenarios. The WebShells were collected from public security repositories and augmented with synthetic obfuscation techniques to increase diversity.  [^3] We acknowledge the potential for data leakage due to the pre-training of LLMs. However, the autoregressive nature of this process does not explicitly capture WebShell classification, although it may affect other code generation tasks. To further minimize any risk of leakage, we restricted our selection of PHP programs to projects updated between October 2024 and 2025, thus reducing overlap with LLM training corpora. Using GPT-4’s tokenizer, we analyzed the token lengths of both sample types. The WebShell samples had a maximum token length of 1,386,438 and an average of 30,856.60 tokens, compared to a maximum of 305,670 tokens and an average of 2,242.89 tokens for benign programs. These results indicate that WebShells are typically significantly longer than benign samples. A detailed summary of the dataset composition can be found in the Table 3 in the Appendix D.
@@ 3.2 Experiment Setup
We randomly selected 60% of the dataset to create a fixed, known demonstration library for ICL. Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing WebShells from benign programs. These scores, detailed in the Table 4 in the Appendix D, guided the selection of ICL demonstrations from the known sample library.
@@
We compared our approach to several baselines: GloVe + SVM, CodeBERT + Random Forest, GCN, and GAT. For GloVe + SVM, we used pre-trained GloVe embeddings with a dimensionality of 300 and an SVM classifier with default parameters (Qi et al., 2018; ZENG et al., 2025). For CodeBERT + Random Forest, we used CodeBERT embeddings with a hidden dimension of 768 and a Random Forest classifier with default settings (Wang et al., 2024a). The GCN was trained with a learning rate of 0.001 over 120 epochs, with 3 hidden layers and a hidden dimension of 32 (Feng et al., 2024). The GAT was trained with a learning rate of 0.001 over 120 epochs, with 3 hidden layers, a hidden dimension of 8, and 8 attention heads (Feng et al., 2024). We evaluated the models using standard classification metrics: accuracy, precision, recall, and F1 score.
@@ 4 Results and Analysis
In this section, we systematically evaluate the performance of LLMs of different scales for WebShell detection and assess the improvements provided by our proposed BFAD framework. Our analysis addresses three research questions (RQs) to explore both the baseline LLM capabilities and the effectiveness of the BFAD components:
@@ 4.1 Performance Evaluation of LLMs and BFAD Enhancements (RQ1)
To answer our first research question, we evaluated seven LLMs against traditional ML and DL baselines. The LLMs included large-scale models (GPT-4, LLaMA-3.1-70B, Qwen-2.5-Coder-14B) and small-scale models (Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, Qwen-2.5-0.5B). Baselines consisted of sequence-based methods (Glove+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT). The detailed performance comparison is presented in Table 1.
@@
Our experimental results yield two core findings. First, vanilla LLMs exhibit a stark, size-dependent trade-off between precision and recall, with all models lagging behind the state-of-the-art GAT baseline. For instance, GPT-4 achieves 100% precision but only 85.98% recall, while the small Qwen-2.5-0.5B model reaches 100% recall but a mere 18.65% precision. This indicates that out-of-the-box LLMs are not well-calibrated for the nuanced task of WebShell detection, making targeted intervention necessary for effective deployment.
Second, the BFAD framework dramatically improves performance across all models, elevating large LLMs to surpass the GAT baseline and making small models viable competitors. With BFAD, GPT-4’s F1 score increases by 6.89 points to 99.35%, exceeding the GAT baseline. The most significant gain is seen in Qwen-2.5-0.5B, whose F1 score improves by a remarkable 51.23 points. This demonstrates that by strategically guiding the model’s focus on behaviorally relevant code, BFAD effectively bridges the performance gap, unlocking the potential of LLMs for this task without resource-intensive retraining.
@@
The performance disparities observed in our results appear to stem from fundamental differences in the models’ analytical reasoning. The precision-recall imbalance in baseline LLMs is a direct reflection of their reasoning strategy. Small models like Qwen-2.5-0.5B tend to rely on surface-level heuristics; they identify keywords like base64_decode and flag them as malicious without assessing the surrounding code’s intent. This leads to high recall but generates numerous false positives when such functions are used legitimately (e.g., in a cryptography library). In contrast, large models like LLaMA-3.1-70B employ deep contextual reasoning, correctly identifying legitimate use cases and thus achieving high precision. However, this sophistication can be a double-edged sword, as these models can be misled by well-disguised WebShells, rationalizing away malicious indicators and causing false negatives that reduce recall. Our proposed BFAD mitigates these issues by guiding smaller models to better understand function-level behaviors and helping larger models focus on critical functions. As a result, it significantly improves detection performance across scales.
@@
We observed a counterintuitive scaling effect where the Qwen-2.5-1.5B model, despite outperforming the 0.5B variant in baseline tests, performed worse with BFAD. We attribute this to differing sensitivities to ICL across scales. The smallest models, like the 0.5B variant, are highly susceptible to label bias and tend to mimic the ICL demonstration’s label without deep reasoning (Fei et al., 2023). Because BFAD provides accurately matched demonstrations via WBFP, the 0.5B model benefits disproportionately from this behavior. In contrast, the 1.5B model begins to exhibit more independent reasoning based on its pre-training knowledge. This emergent but still imperfect reasoning can conflict with the ICL example’s guidance, degrading performance. This aligns with prior findings that LLMs sometimes rely more on pre-training than on provided examples (Peng et al., 2025). Therefore, the 0.5B model’s superior performance with BFAD does not indicate greater intelligence but is an artifact of its strong alignment with the high-quality examples provided.
@@ 4.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
We evaluated the effectiveness of Context-Aware Code Extraction using two models: GPT-4, a large-scale model, and Qwen-2.5-3B, a smaller-scale model. Three configurations were compared: (1) predictions based on the full source code, (2) predictions using only extracted critical regions, and (3) a hybrid approach combining critical regions with truncated source code. Results are reported in Tables 5 and 6 in Appendix E.
@@
For the smaller model, Qwen-2.5-3B, critical regions significantly improve performance over the full source baseline. At $\tau=100$, the F1 score increases from 84.37% to 90.91% (+6.54%), with precision increasing from 78.03% to 86.71% (+8.68%) and recall increasing from 91.84% to 95.54% (+3.70%). This improvement validates the focus on behaviorally relevant code snippets, which reduces irrelevant context and sharpens the focus of the model. However, as the context length ($\tau$) increases, performance decreases - F1 drops to 88.17% at $\tau=300$, likely due to the model’s limited ability to handle extended context, which introduces noise that degrades accuracy.
For the larger model, GPT-4, critical regions increase recall but slightly decrease precision. At $\tau=300$, recall improves from 85.98% to 96.18% (+10.20%), while precision drops from 100.00% to 98.69% (-1.31%). The F1 score increases from 92.46% to 97.42% (+4.96%), indicating that GPT-4 effectively uses localized behavioral cues to improve recall. However, the reduced global context may introduce small biases, leading to a precision trade-off, although the overall performance remains strong due to the model’s greater capacity.
@@
The hybrid strategy improves model performance by effectively balancing precision and recall. For Qwen-2.5-3B, it increases precision over using critical regions alone, from 86.71% to 89.02% (+2.31%) for $\tau=100$ and from 82.32% to 85.55% (+3.23%) for $\tau=300$. Although recall decreases slightly, the F1 score increases to 89.70% at $\tau=300$ (+1.53% from 88.17%), indicating that the hybrid approach reduces noise in longer contexts and supports smaller models by maintaining focus on critical regions while integrating valuable global context. For GPT-4, the strategy improves recall without compromising precision: at $\tau=300$, recall increases from 85.98% (using the source code) to 96.82% (+10.84%), while precision remains steady at 100.00%, yielding an F1 score of 98.38%–a 5.92% gain over the source code and a 0.96% improvement over critical regions alone.
@@ 4.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4. This evaluation builds on the optimal hybrid strategy from RQ2. Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt). Results are detailed in Tables 7 and 8 in Appendix E.
@@
Random selection makes ICL performance much worse by adding irrelevant examples, which makes both models much less effective. For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 30.14% reduction from the no-ICL baseline of 90.97%. This is mainly because there is a large drop in precision from 89.02% to 46.33%, although recall remains strong at 88.53%. Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 22.16% decrease from the No-ICL baseline of 98.38%, driven by a drop in precision from 100.00% to 65.58%. These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
The SC-Sim approach, which relies on semantic similarity calculated over entire source code samples, also doesn’t work well. For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, a notable decrease from the No-ICL baseline of 90.97%, with a precision of 78.57% and a recall of 91.08%. For GPT-4, it achieves an F1 score of 96.32%, with perfect precision, but a reduced recall of 92.90% compared to the No-ICL baseline’s 96.82%. This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.
@@
For both models, WBFP-Wt consistently outperforms in accuracy, precision, recall, and F1 score, demonstrating its robustness and adaptability for improving ICL in WebShell detection tasks. Specifically, for Qwen-2.5-3B, WBFP-Wt achieves an F1 score of 93.69%, which is 9.33% higher than SC-Sim and 1.28% higher than WBFP-Eq. This is because WBFP-Wt achieves a precision of 88.64% (compared to 78.57% for SC-Sim and 86.39% for WBFP-Eq) and a near-perfect recall of 99.36%. By focusing on the important parts of WebShell behavior, WBFP-Wt makes up for the fact that the smaller model doesn’t understand as much. This results in demonstrations that closely match the desired behavioral profiles, improving both precision and recall.
For GPT-4, WBFP-Wt achieves the highest F1 score of 99.35%, which is 3.03% higher than SC-Sim and 0.34% higher than WBFP-Eq . While precision remains at 100.00% across all WBFP variants, WBFP-Wt increases recall to 98.71%, compared to 92.90% for SC-Sim and 98.03% for WBFP-Eq. This improvement highlights WBFP-Wt’s ability to leverage GPT-4’s advanced understanding of context, matching selected demonstrations to the behavioral characteristics of the target sample to optimize recall without sacrificing precision.
5 Related Work
WebShell Detection Techniques Early efforts in WebShell detection relied on rule-based methods that used signature matching or heuristics to identify malicious scripts (Le et al., 2021; Jinping et al., 2020). For example, Le et al. (2021) proposed H-DLPMWD, a hybrid approach that combines pattern matching with a CNN to detect ASP.NET WebShells, achieving 98.49% accuracy by using Yara-based filtering and opcode indexing. However, such methods struggle against obfuscated or novel WebShell variants due to their reliance on predefined patterns (Hannousse & Yahiouche, 2021). Machine learning (ML) techniques have advanced this landscape by extracting features from code text or runtime behavior. Jinping et al. (2020) introduced a mixed-model approach using Random Forest and CNNs with N-gram and TF-IDF features that achieved 97% accuracy on PHP WebShells, but it requires balanced datasets and struggles with encrypted samples. Deep learning (DL) has further improved adaptability, with models such as CodeBERT for semantic analysis (Pu et al., 2022) and Graph Attention Networks (GAT) for structural insights (Zhang et al., 2025). (Zhang et al., 2025) proposed MMFDetect, which fuses CodeBERT-CL semantics with CNN-extracted visual features from RGB-mapped PHP code, achieving 99.47% accuracy on evasive WebShells. Despite these gains, ML and DL methods require extensive labeled data-a scarcity in cybersecurity-and exhibit limited generalization to obfuscated threats, along with high computational cost (Shang et al., 2024; Jinping et al., 2020).
LLMs in Code Analysis LLMs have revolutionized code-related tasks by leveraging large pre-training corpora for applications such as code generation (Ma et al., 2024), vulnerability detection (Sun et al., 2024), and program reliability assessment (Liu et al., 2024). Ma et al. (2024) demonstrated the ability of LLMs to generate evasive WebShells using hybrid prompts, highlighting their code synthesis potential. Sun et al. (2024) introduced LLM4Vuln, which enhances vulnerability reasoning through knowledge retrieval and has achieved practical success in Solidity audits. These models excel at zero-shot and few-shot learning via prompt engineering (Nong et al., 2024), providing interpretability critical for security contexts (Ma et al., 2024). However, their application to WebShell detection remains underexplored, with previous studies focusing on generation or general vulnerabilities rather than detection of stealthy, context-heavy WebShells.
Challenges of LLMs for WebShell Detection Using LLMs for WebShell detection reveals critical bottlenecks. The fixed context window truncates large WebShells, potentially missing malicious segments embedded in benign code (Ceka et al., 2024; Wang et al., 2025). Fang et al. (2024) found that LLM performance degrades with longer inputs, with GPT-4’s accuracy dropping to 87% on obfuscated JavaScript. Solutions such as chunking or sparse attention (e.g., LongCoder (Guo et al., 2023), SparseCoder (Wang et al., 2024b)) mitigate this, but often lose global context (Wang et al., 2025). In-context learning (ICL), a cornerstone of LLM adaptability, falters because demonstrations consume context space (Min et al., 2022; Wang et al., 2025), and random or semantic similarity-based selections fail to capture WebShell-specific behaviors. Liu & Wang (2023) proposed maximum information gain for ICL, but its focus on text classification limits its applicability to code security.
These gaps between rule-based rigidity, ML and DL data dependency, and LLM context and ICL limitations motivate our BFAD framework. BFAD overcomes context constraints with a hybrid extraction strategy that preserves critical regions and global cues, outperforming naive LLM applications. In addition, our WBFP enhances ICL by prioritizing behaviorally relevant demonstrations, outperforming generic similarity-based selection methods.

### v5

@@ 3.4 LLM-Based Detection Framework
We integrate BFAD into an LLM-based detection pipeline that combines context-aware extraction with behavior-weighted demonstration selection. By providing behavior-relevant snippets together with complementary global context and a closely matched demonstration, the LLM is better positioned to distinguish malicious intent from benign uses of similar APIs.
The LLM input consists of (a) a system directive specifying the cybersecurity-analysis role and (b) a user query containing the extracted code view and the selected ICL demonstration. To balance effectiveness and cost, we include a single demonstration per query. The full prompt template is provided in Appendix A.
@@ 4.1 Dataset Overview
We construct a dataset of 26,594 PHP scripts, including 21,665 benign programs and 4,929 WebShells. Benign programs are sourced from established open-source PHP projects to reflect realistic development practices. WebShells are collected from public security repositories and further augmented with synthetic obfuscation to increase behavioral and lexical diversity. [^2] We acknowledge the risk of data leakage due to LLM pre-training. While pre-training does not explicitly optimize for WebShell classification, memorization of code snippets could still affect evaluation. To reduce overlap with potential training corpora, we restrict benign projects to those updated between October 2024 and 2025. To characterize the distributional shift and the challenge posed by long contexts, we measure sequence lengths using the GPT-4 tokenizer. WebShells are substantially longer than benign files, with a maximum of 1,386,438 tokens and an average of 30,856.60 tokens, compared to a maximum of 305,670 and an average of 2,242.89 tokens for benign programs. A detailed breakdown of the dataset is provided in Table 3 (Appendix C).
@@ 4.2 Experiment Setup
We randomly sample 60% of the dataset to form a fixed demonstration library for ICL. Using this subset, we compute WBFP-based weights for each critical function type by combining coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$). The resulting normalized weights quantify each function type’s discriminative importance for separating WebShells from benign programs and are used to rank candidate demonstrations. The full weights are reported in Table 4 (Appendix C).
@@
We compare BFAD-enhanced LLMs against representative classical and graph-based detectors: GloVe+SVM, CodeBERT+Random Forest, GCN, and GAT. For GloVe+SVM, we use 300-dimensional pre-trained GloVe embeddings and an SVM classifier with default hyperparameters (Qi et al., 2018; ZENG et al., 2025). For CodeBERT+Random Forest, we use CodeBERT embeddings (768 dimensions) with a Random Forest classifier using default settings (Wang et al., 2024a). For the graph models, we follow prior work (Feng et al., 2024): both GCN and GAT are trained for 120 epochs with learning rate 0.001; GCN uses 3 hidden layers with hidden size 32, and GAT uses 3 hidden layers with hidden size 8 and 8 attention heads. We report standard classification metrics: accuracy, precision, recall, and F1.
@@ 5 Results and Analysis
This section evaluates LLMs of different scales for WebShell detection and quantifies the gains delivered by our proposed BFAD framework. We structure the analysis around three research questions (RQs), covering both out-of-the-box LLM behavior and the contribution of individual BFAD components:
@@ 5.1 Performance Evaluation of LLMs and BFAD Enhancements (RQ1)
To address RQ1, we compare seven LLMs with representative classical detectors. The LLMs include large-scale models (GPT-4, LLaMA-3.1-70B, Qwen-2.5-Coder-14B) and smaller models (Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, Qwen-2.5-0.5B). Baselines include sequence-based methods (GloVe+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT). Table 1 reports all metrics.
@@
Our results highlight two key observations. First, vanilla LLM prompting exhibits a pronounced, scale-dependent precision–recall trade-off and remains below the strongest graph baseline (GAT). For example, GPT-4 attains 100% precision but only 85.98% recall, whereas Qwen-2.5-0.5B achieves 100% recall with only 18.65% precision. This asymmetry suggests that off-the-shelf LLMs are poorly calibrated for WebShell detection, motivating behavior-aware intervention.
Second, BFAD improves every evaluated LLM, elevating large models beyond the GAT baseline and making small models substantially more competitive. With BFAD, GPT-4 improves by +6.89 F1 to 99.35%, surpassing GAT. The largest relative gain occurs for Qwen-2.5-0.5B (+51.23 F1), indicating that explicitly steering attention toward behavior-relevant code and demonstrations can unlock strong performance without task-specific fine-tuning.
@@
The above gap is consistent with qualitative misclassification patterns. Smaller models (e.g., Qwen-2.5-0.5B) often rely on surface heuristics: they over-trigger on high-risk API names such as base64_decode and ignore surrounding intent, yielding high recall but many false positives when such APIs appear in benign utilities (e.g., cryptography libraries). In contrast, larger models (e.g., LLaMA-3.1-70B) more frequently leverage broader context and therefore maintain high precision, but can be distracted by plausible-looking surrounding logic and miss carefully disguised WebShells, producing false negatives. BFAD mitigates both failure modes by anchoring analysis on behavior-critical calls while still preserving enough local (and partial global) context for intent disambiguation.
@@
We observe a counterintuitive scaling pattern: Qwen-2.5-1.5B, while stronger than the 0.5B variant under vanilla prompting, improves less under BFAD and remains worse than Qwen-2.5-0.5B with BFAD. A plausible explanation is scale-dependent sensitivity to ICL. Very small models are more prone to label bias and may over-imitate the demonstration label rather than perform independent reasoning (Fei et al., 2023); when WBFP retrieves well-matched demonstrations, this bias can translate into large gains. As model size increases, the model may rely more on pre-training priors and partially resist demonstration guidance, which can introduce conflicts and reduce the net benefit (Peng et al., 2025). Thus, the strong BFAD performance of the 0.5B model should be interpreted as a demonstration-following effect rather than superior underlying reasoning.
@@ 5.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
To answer RQ2, we evaluate Context-Aware Code Extraction on a large model (GPT-4) and a smaller model (Qwen-2.5-3B). We compare three input configurations: (1) the full source code, (2) extracted critical regions only, and (3) a hybrid input that concatenates critical regions with additional truncated code segments. Results are reported in Tables 5 and 6 (Appendix D).
@@
For Qwen-2.5-3B, focusing on critical regions substantially improves performance over full-code prompting. At $\tau=100$, F1 increases from 84.37% to 90.91% (+6.54), with precision increasing from 78.03% to 86.71% (+8.68) and recall increasing from 91.84% to 95.54% (+3.70). This supports the hypothesis that removing irrelevant context sharpens behavioral evidence for smaller models. However, increasing $\tau$ can re-introduce noise: at $\tau=300$, F1 drops to 88.17%, consistent with limited robustness to longer, noisier contexts.
For GPT-4, critical regions mainly improve recall with a minor precision trade-off. At $\tau=300$, recall rises from 85.98% to 96.18% (+10.20), while precision decreases from 100.00% to 98.69% (-1.31), resulting in an F1 increase from 92.46% to 97.42% (+4.96). This indicates that localized behavioral cues are particularly helpful for recovering missed attacks, while removing global context can slightly increase false positives.
@@
The hybrid strategy improves the precision–recall balance by re-introducing limited global context. For Qwen-2.5-3B, it increases precision relative to critical-regions-only (86.71% $\rightarrow$ 89.02% at $\tau=100$; 82.32% $\rightarrow$ 85.55% at $\tau=300$). Although recall decreases slightly, the hybrid input yields a higher F1 at $\tau=300$ (89.70%, +1.53 over 88.17), suggesting that carefully chosen global snippets reduce false positives when longer windows add noise. For GPT-4, the hybrid input improves recall without sacrificing precision: at $\tau=300$, recall increases from 85.98% (full code) to 96.82% (+10.84) while precision remains 100.00%, yielding an F1 of 98.38% (+5.92 over full code and +0.96 over critical-regions-only).
@@ 5.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
To answer RQ3, we evaluate WBFP for ICL demonstration selection on Qwen-2.5-3B and GPT-4, using the best-performing hybrid input from RQ2. We compare four retrieval strategies: Random selection, source-code semantic similarity (SC-Sim), WBFP with equal weights (WBFP-Eq), and WBFP with function-level weights (WBFP-Wt). Results are reported in Tables 7 and 8 (Appendix D).
@@
Random demonstrations substantially degrade ICL performance, indicating that irrelevant examples introduce strong prompt noise. For Qwen-2.5-3B ($\tau=100$), F1 drops to 60.83% (from 90.97% without ICL), driven primarily by a precision collapse (89.02% $\rightarrow$ 46.33%) while recall remains high (88.53%). For GPT-4 ($\tau=300$), F1 decreases to 76.22% (from 98.38%), largely due to reduced precision (100.00% $\rightarrow$ 65.58%). These results show that naive ICL can be harmful without behaviorally relevant retrieval.
SC-Sim, which computes semantic similarity over entire source files, is also suboptimal. For Qwen-2.5-3B, SC-Sim yields 84.36% F1 (vs. 90.97% without ICL), with precision 78.57% and recall 91.08%. For GPT-4, SC-Sim reaches 96.32% F1 with perfect precision but reduced recall (92.90% vs. 96.82%). We hypothesize this occurs because whole-file similarity is dominated by behavior-irrelevant regions, which dilutes the behavioral cues needed for effective WebShell retrieval.
@@
WBFP improves ICL reliably, and WBFP-Wt is consistently the strongest variant. For Qwen-2.5-3B, WBFP-Wt achieves 93.69% F1, improving over SC-Sim by +9.33 and over WBFP-Eq by +1.28. The gain comes from both higher precision (88.64% vs. 78.57% for SC-Sim) and near-perfect recall (99.36%). These results suggest that behavior-weighted, function-centric retrieval is particularly beneficial for smaller models whose general code reasoning is limited.
For GPT-4, WBFP-Wt achieves 99.35% F1 (+3.03 over SC-Sim and +0.34 over WBFP-Eq). Precision remains 100.00% across WBFP variants, while WBFP-Wt improves recall to 98.71% (vs. 92.90% for SC-Sim and 98.03% for WBFP-Eq). This indicates that behavior-aware retrieval primarily helps large models recover hard positives without sacrificing precision.
6 Limitations and Future Work
In this section, we outline several concrete directions for future work—including synthetic benchmark construction, graph–LLM integration, and agentic detection pipelines—with the goal of making our ideas actionable and easy to build upon.
6.1 Dataset Generalization
A central challenge in WebShell research is the limited availability of diverse, real-world data. As in prior work, we primarily rely on public datasets, which may not fully reflect emerging, targeted, or highly obfuscated threats, and therefore limits our ability to rigorously assess generalization. An important direction is to use LLMs to construct large-scale synthetic benchmarks that are both privacy-preserving and diverse, enabling stress tests across a broader spectrum of attack vectors.
6.2 Model and System Robustness
Our framework is primarily static, and thus remains susceptible to advanced obfuscation and previously unseen (e.g., zero-day) behaviors. Future work should improve both robustness and adaptability. To mitigate evasive variants, one could incorporate dynamic analysis, e.g., by extracting runtime traces from sandboxed execution (Zhao et al., 2024; Han, 2026).
To strengthen semantic understanding, domain-specific fine-tuning of code-oriented LLMs on curated WebShell corpora is another promising direction. Beyond standard fine-tuning, we propose a concrete next step that frames WebShell detection as a multimodal fusion problem: prior work suggests that graph representations can capture program behavior (e.g., function-call structure) more faithfully than pure sequence models (Han, 2026). Building on this observation, future work could (i) train a dedicated graph encoder to map code into behavior-centric embeddings (instead of relying solely on the LLM’s text embeddings), and then (ii) learn an adapter that aligns the graph embeddings with natural-language instructions and the LLM’s latent space (in the spirit of adapter-based alignment in VLMs and emerging graph-LLM systems). We believe this design can combine the strengths of both worlds: robust behavioral feature extraction from graphs, and the LLM’s pretrained discrimination and interpretability.
For deployment, a practical direction is a multi-model agentic system that orchestrates specialized components. These components naturally fit a multi-stage “fast–slow” pipeline: a lightweight model performs high-recall triage, followed by a larger model that conducts deeper analysis on flagged samples. Finally, to sustain performance under distribution shift, the pipeline could be augmented with an autonomous update loop that discovers new patterns and refreshes detection models over time.